How much are the wheat rolls and how much are the rye rolls.
4 Wheat rolls + 4 Rye rolls = $4
6 wheat rolls + 2 rye rolls = $3.60

Answer by TimothyLamb(4379)

x = dollar price per wheat roll
y = dollar price per rye roll
---
4x + 4y = 4
6x + 2y = 3.60
---
put the system of linear equations into standard form
---
4x + 4y = 4
6x + 2y = 3.60
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 0.40
y = 0.60
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---